Average undergraduate cost for tuition, other fees, and boarding for all institutions in the US last year is $19410. A random sample of costs this year for 40 institutions of higher learning indicated that the sample mean was $22098, and the sample standard deviation was $6050. At 0.01 level of significance, is there sufficient evidence to conclude that the cost has increased


1
Expert's answer
2021-10-18T15:07:02-0400

Solution:

We are given a sample standard deviation but since the sample size is greater than 30, we can use the z distribution instead of the t distribution.

"\begin{aligned}

&H_{0}: \mu=19410 \\

&H_{1}: \mu>19410

\end{aligned}"  

Test Statistic: "z=\dfrac{\bar{x}-\mu}{\sigma / \sqrt{n}}"

Rejection Region: "\left\{z: z \geq z_{\alpha}=z_{0.01}=2.33\right\}"

 "z=\dfrac{22098-19410}{6050 / \sqrt{40}}=2.81"  

Reject H0 since the value of the test statistic 2.81, is greater than the critical value 2.33.

Thus, there is sufficient evidence to conclude that the cost of attendance has increased